OpenStudy (anonymous):

Find the equation of the line that passes through (4, 5) with a slope of 2. Write your answer in Slope-Intercept Form.

OpenStudy (anonymous):

y - 5 = 2(x - 4)

OpenStudy (katrinakaif):

Plug it into the equation: y-y1 = m (x-x1) with given values of m = 2 and (4,5) : (x1,y1)

OpenStudy (anonymous):

Thank you Aries and Katrina. (:!!!

OpenStudy (katrinakaif):

The pleasure is mine :-)

OpenStudy (anonymous):

y - 5 = 2x - 8 y = 2x + 3 = SLOPE INTERCEPT

OpenStudy (anonymous):

u r welcome;)